OpenStudy (anonymous):

I'm confused on functions...if f(x)=2x+1, what is f(t+1) ?

OpenStudy (usukidoll):

f(x) = f(t+1) it means that x = t+1

OpenStudy (usukidoll):

so for your function f(x) = 2x+1 and we're asked f(t+1). that means that our x = t+1. So just plug that into our function f(t+1) = 2(t+1)+1 and solve

OpenStudy (anonymous):

ohhh so is it okay to still have a variable, in this case t, in the solution?

OpenStudy (usukidoll):

yup

OpenStudy (usukidoll):

so just solve this part 2(t+1)+1 and that will be our f(t+1)

OpenStudy (usukidoll):

distribute the 2 for (t+1) 2(t)+2(1)+1 and combine like terms/simplify

OpenStudy (usukidoll):

what's 2+1 ?

OpenStudy (anonymous):

=2t+3 ?

OpenStudy (usukidoll):

correct

OpenStudy (anonymous):

thank you so much!!

OpenStudy (usukidoll):

THANK YOU! This question got me to my goal of 90SS!!!!!!!!!!